|
|
|
IRIS2 Goes to the AAT |
|
|---|---|
|
|
|
John Dawson's Diary |
|
|
|
|
October 23 |
|
| Arrived yesterday at around 15:30 after a drive through considerable
rain. The truck with IRIS2 and Roger on board arrived within the hour.
We had the truck off loaded by about 18:00. Moved all the equipment upto
the 4th floor and imediately set about checking the fit of IRIS2 to the
cass. Offered the whole instrument up and it fitted perfectly (good one
Greg).
Then set about getting main dewar pumped down and leak checked. While this was in progress the helium lines were setup in preparation for the cool down. This lead to trying IRIS2 in a number of parked position as various interface problems came to light (not major just reach of cables etc.). We are now parked in the 4th position since we arrived. The main dewar was found to have a small leak and we worked until 2:00 this morning to try and rectify it. We were not successful and we are continuing with the very tiny leak. John |
|
|
|
|
| Today we have been leak checking the fore dewar and this
has also shown a leak at the pressure transducer, of a larger magnitude
than the one found in the main dewar. We are still trying to rectify this
leak and are confident we will have it fixed in the next few hours. We
have not been rushing as we were waiting for the appropriate software to
be available to run the wheels warm before commencement of cooling. With
Rolf busying himself with all the cabling today they have probably run
the wheels while I have been writing this email.
Jeremy and Don arrived around 15:30 and Don has already mounted the controller on the instrument. Before today is out we believe that we will have both dewars pumped and cooling. Tomorow we will offer the instrument back up to the telescope to put in the final structural pins. The instrument will be in position tomorow night but will probably be to warm to do any meaningful observing. However, this will be in the hands of Chris and Roger. As I understand it they are confident they will be at temperature and on the sky by Saturday night. John |
|
October 25 2001IRIS2 was cooled overnight without incident. It has been put on the telescope this morning without any problems. The helium lines have been connected and the cryocoolers are running via the site installed helium lines. The fore dewar is still on the vacuum pump to take its pressure down to an acceptable level.Denis and some of the site team are now putting in the remaining pins to complete the external structure. At the moment I think Chris might try and take images with IRIS2 tonight, but that remains his call. All being well most of the Epping team will be attending 'Science in the Pub' tonight in Coona. A well deserved break I believe. John. |
|
![]()
This page created by Chris Tinney